OpenStudy (anonymous):

How do i solve (x^3+64)

OpenStudy (lgbasallote):

=(x^3 + 4^3) =(x+4)(x^2 -4x + 16)

OpenStudy (anonymous):

can someone explain how to do it please

OpenStudy (lgbasallote):

there is a principle.. (a^3+ b^3) = (a+b)(a^2 -ab + b^2)

OpenStudy (anonymous):

solve or factor? u need an = something or an x value to solve

OpenStudy (anonymous):

factor, sorry

OpenStudy (lgbasallote):

one thing you should also know since it might be useful is (a^3-b^3) = (a-b)(a^2 + ab + b^2)